OpenStudy (anonymous):

Express as a single log: 1/2log5 x+2log5(x+2)-[log5(1+x)+log5(1-x)]

OpenStudy (anonymous):

When you say log5 do you mean that 5 is the base of the log?

OpenStudy (anonymous):

Here are the four essential log identities: \[\log a + \log b = \log (ab)\] \[\log a - \log b = \log (a/b)\] \[a \log b = \log b ^{a}\] \[\log _{a}b = (\log b) / (\log a)\] The second two follow from the first one. All log problems are just applying one of these rules at the right time.